Starting from Rest, a block of mass m = 5.0kg slides down a frictionless ramp. The ramp has a vertical height of H= 1.0 meter and makes an angle of 30 degrees with the floor. After leaving the ramp, the block slides along the floor. But the floor is not frictionless. In fact, the floor'S coefficient of kinetic friction is 0.20.


A) What is the block's speed at point B? Solve this without any kinetic formulas


B.) How much work does the floor do on the block as the block slows down and eventually comes to rest?


C.) How far along the floor does the block slide before coming to rest?


D.) How would your answer to parts (b) and (c) change if the floor's coefficient or friction were =. 40 instead of =. 20? Just answer in qualitatively how things would be different.


1
Expert's answer
2021-06-01T10:50:17-0400

a)


"v^2=2(9.8)1\\v=4.4\frac{m}{s}"

b)


"W=mgh=(5)(9.8)(1)=49\ J"

c)


"49=0.2(5)(9.8)l\\l=5\ m"

d) The work does not change. The length will be the half of initial length.